# Benchmarking General Purpose In-Context Learning

### v2

@@ 3.2 Baseline Solution
We offer a baseline solution for the Maze World tasks, which is largely inspired by model-based imitation learning (Hu et al., 2022), causal decision models (Chen et al., 2021), and meta-reinforcement learning (Duan et al., 2016). We use Variational Auto Encoder (VAE) to process each image to latent representation, and use a causal transformer (Chen et al., 2021) to encode the historical observation and action. The model ouputs the distribution of actions, and also the expected next frame. Specifically, given the reference trajectory $(i_{1},a_{1},i_{2},a_{2},...,i_{T},a_{T},i_{T+1})$, where $i_{t}$ and $a_{t}$ represent the observation and action respectively, our model is represented by:
@@
We train the model using imitation learning, although reinforcement learning is desirable for future enhancements. The losses are composed of three parts: 1. the reconstruction error ($\mathcal{L}_{vae}$); 2. the cross entropy between the reference action and the predicted distribution ($\mathcal{L}_{pm}$); 3. the next-frame prediction error ($\mathcal{L}_{wm}$). The final loss would be:
@@
$\mathcal{L}$ $=\alpha_{1}\mathcal{L}_{vae}+\alpha_{2}\mathcal{L}_{pm}+\alpha_{3}\mathcal{L}_{wm}$ (8)
The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (330M parameters). A overview of the model architecture is in Fig. 7. To train our baseline models, we collect training data include 12.8K episodes of demonstration from rule agents for imitation learning. To avoid the growing “compounding errors” (Ross and Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from data augmentation methods (Ross et al., 2011). We collected the trajectory of states by running a very noisy behavior policy, but used a rule-based agent to label its action at each step (which is not actually executed; see appendices for details). Each episode has $2,048$ steps, which gives $26$M frames in total.
Figure 7: An overview of the model evaluated on the maze world tasks. Initially, a Variational Auto-Encoder (VAE) is trained to encode observed images into hidden vectors and vice versa, with its parameters subsequently frozen. The sequence of observations and actions is then utilized to train a causal model, specifically a transformer that employs only backward attention, using imitation learning and self-supervised learning techniques.
@@ 3.3 Baseline Performances
Figure 8: Static assessment of the ability to mimic rule-based agents ($\mathcal{L}_{pm}$) and predict the next frame ($\mathcal{L}_{wm}$) across each time step ($t$). Context lengths of 512 and 2048 are evaluated respectively for comparison.
Static Evaluation. We first conduct static evaluation using off-policy data from an independently sampled validation set comprising $720$ sequences (totaling $1.5$M frames). To analyze the impact of context length on performance, we assessed the losses of the world model and policy model for each individual time step ($\mathcal{L}_{wm}(t)$ and $\mathcal{L}_{pm}(t)$). In Figure 8, We compare the performance of the small-sized and standard-sized causal models, with several noteworthy observations: First, the future prediction task shows no clear dependency on long contexts (the learning horizon is below 100 steps), whereas the performance of action prediction continually improves with longer contexts. This likely suggests that world modeling in the proposed tasks is more related to few-shot ICL capabilities than to GPICL. This is unsurprising, since single-step world modeling is more connected with understanding the localized environment, whereas navigation requires more global reasoning. Second, the standard-sized model does not demonstrate a clear advantage over the small-sized model. This phenomenon aligns with that observed in meta-language tasks, further validating the idea that scaling up parameters is not always necessary compared to scaling up contexts in GPICL.
@@
Figure 9: Interactive evaluation results of the rule-based agents, randomized policy, and the causal model, depicted by plotting accumulated rewards against simulated time steps. Each plot summarizes the results from identical 32 tasks. The step minor cost is set to 0.
Interactive Evaluation. It is more beneficial to investigate the online performance of the policy model and world model in causal modeling by directly interacting with the environment.
@@
We first directly evaluate the model-free policy by interacting with a set of evaluation tasks. The evaluation set includes 32 tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$. A softmax policy, based on the predicted distribution of actions, is implemented. We plot the mean and variance of the average accumulated rewards over time (t). For comparative analysis, we also include the performance of a random policy, where actions are randomly sampled regardless of the observations, and agents with different short-term memory to long-term memory transition probabilities $p(\text{STM}\rightarrow\text{LTM})$ of $5\%$, $25\%$, $100\%$ (Figure 9). The results indicate that our model significantly outperforms random exploration and rule-based agents with low long-term memory capabilities ($p(\text{STM}\rightarrow\text{LTM})=5\%$). However, there is a noticeable gap compared to those with more effective long-term memories ($p(\text{STM}\rightarrow\text{LTM})=25\%,100\%$). Moreover, the rate of improvement is not significant compared to those with effective long-term memories, where we would expect to see an accelerating accumulation of rewards. We believe that there is ample room for improvement by employing more sophisticated models and better learning algorithms, and extending the context length.
Figure 10: Randomly selected clips depicting the future 8 steps using the world model of causal modeling are presented. For each clip, the upper section shows the ground truth observations from the current step to 8 steps ahead (from left to right), while the bottom section shows the imagined future up to 8 steps ahead starting from the current observation. The actions taken are also labeled with text in the image. The future imagination is generally accurate for the first frame but deteriorates over time due to error propagation, illustrating the challenge of causal world modeling.
We then provide a qualitative evaluation of the performance of world models in Figure 10. Unlike $\mathcal{L}_{wm}$, which can only represent the precision of predicting one-step transitions, using the world model to predict the distant future is far more challenging. We observe that although our world model can predict the nearest future, its accuracy quickly deteriorates over time. This can be attributed to the well-known problem of horizon-square compounding errors. Typical improvements include adding noise to the inputs (Hu et al., 2022) and multi-step prediction (Asadi et al., 2019). However, compared to many previous problems, Maze World poses a higher challenge to world modeling in two key aspects: First, the variance in observations is more severe (the observation of two consecutive frames can be very different). Second, the variance of the tasks requires the world model to perform in-context learning to improve, as each maze is completely different and not transferrable.
@@ 4.1 Meta-learning and GPICL
General purpose in-context learning (GPICL) is closely related to meta-learning , particularly model-based meta-learning (Duan et al., 2016; Santoro et al., 2016). The latter generally encompasses a wide range of methodologies, including those that rely on gradient-based adaptation (Finn et al., 2017), while in-context learning focuses on gradient-free “instruction following” (Ouyang et al., 2022). In recent years, the distinction between meta-learning and pre-training-ICL has become increasingly blurred (Reed et al., 2022; Dai et al., 2023; Li et al., 2023). However, GPICL fundamentally differs from traditional model-based meta-learning and ICL in large-scale language models in the key aspects including minimal zero-shot generalization, long learning horizon, and abundancy of learning algorithm (where supervised learning, reinforcement learning and others are included).
4.2 Benchmarks for Meta-Learning and ICL
A variety of benchmarks have been utilized to evaluate meta-learning algorithms, encompassing domains such as image classification (LeCun et al., 1998), video games, and robot manipulation (Brockman et al., 2016). Many of these benchmarks were not initially designed for meta-learning but have been adapted to accommodate varying tasks through parametric adjustments in hyper-parameters, such as target positions and robot configurations. Meta-World (Yu et al., 2020) introduces a series of robot arm manipulation tasks that incorporate both parametric and non-parametric variations. Mazes are also frequently used to benchmark meta-reinforcement learning (Duan et al., 2016; Mishra et al., 2018), but the scale and diversity of those tasks are not adequate enough to benchmark GPICL.
A key distinction between our proposed benchmarks and previous ones in meta-learning is that the latter often veer towards "task identification" by selectively omitting critical observations. For example, scenarios involve a robot moving in a direction that remains hidden and can only be inferred from rewards (Finn et al., 2017), or a robot arm reaching for a target that is neither visible nor directly observed (Yu et al., 2020). In practical settings, such information like targets and directions would typically be explicitly indicated through instructions or visual cues, making the tasks ostensibly complex yet fundamentally simplistic. Moreover, these tasks generally do not necessitate long-term interaction, with most being solvable through relatively brief explorations. Our benchmarks, in contrast, align more closely with real-world challenges. For instance, the Maze World “navigation” tasks can mirror the adaptive learning process of a household robot adapting to a new indoor scenario, where it needs to simultaneously solve localization, mapping, and planning. Another group of Maze World “survival” tasks simulate an agent exploring a new environment to learn about helpful and harmful elements.
Another set of benchmarks has emerged alongside large language models (LLMs), focusing on evaluating the capability of modeling long sequences. The Long Range Arena (Tay et al., 2020) includes six tasks, primarily adapting image classification tasks to pixel-wise classification. Other benchmarks involve natural corpora, such as literary texts (Rae et al., 2019) and long document classification (Chalkidis et al., 2019). However, a significant issue with using these benchmarks to study GPICL is that it is often difficult to distinguish between GPICL capabilities and zero-shot generalization. Indeed, previous research on efficient transformers (Tay et al., 2022) has shown that improvements in zero-shot generalization (such as through simply increasing model scale) tend to outperform increases in context length (especially beyond 4,000 tokens) in these datasets. This suggests that these benchmarks may not be ideal for thoroughly testing GPICL.
@@ 5 Conclusions
In this paper, we enrich and redefine the topic of General Purpose In-Context Learning (GPICL) and underscore its significance to the AI community. We introduce two benchmarks specifically designed for GPICL. Our preliminary investigations of these benchmarks have yielded some inspiring results, demonstrating that a completely randomized meta-language can be used to train a language model capable of understanding realistic natural language to a certain extent.
It is important to note that although we are seeing the potential of application of those datasets in areas including language modeling and indoor navigation, the synthetic data generated by those benchmarks are not yet validated to be ready for a direct help to application. Currently, large language models (LLMs) typically feature vocabulary sizes $N$ exceeding tens of thousands, highlighting that the true complexity of natural language far surpasses $8$-gram. Moreover, the visual signals in maze worlds remain regularized and relatively monotonic, still significantly less complex than those found in the real world.
@@
This work can be expanded along two directions: 1. The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2. Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.

### v5

@@ 3.2 Baseline Solution
The baseline solution for the Maze World tasks is inspired by model-based imitation learning (Hu et al., 2022) and causal decision models (Chen et al., 2021). Although running meta-reinforcement learning (MetaRL) could be beneficial, we did not directly apply RL due to the significant challenges in scaling RL to meet the required needs. Following previous work that suggests imitating an expert as adequate to generate reinforcement learning capabilities (Lin et al., 2023; Zisman et al., 2023; Lee et al., 2024), we mainly focus on imitation learning augmented with world models. We use Variational Auto Encoder (VAE) to process each image to latent representation, and use a causal transformer (Chen et al., 2021) to encode the historical observation and action. The model ouputs the distribution of actions, and also the expected next frame. Specifically, given the reference trajectory $(i_{1},a_{1},i_{2},a_{2},...,i_{T},a_{T},i_{T+1})$, where $i_{t}$ and $a_{t}$ represent the observation and action respectively, our model is represented by:
@@
The learning losses are composed of three parts: 1. the reconstruction error ($\mathcal{L}_{vae}$); 2. the cross entropy between the reference action and the predicted distribution ($\mathcal{L}_{pm}$); 3. the next-frame prediction error ($\mathcal{L}_{wm}$). The final loss would be:
@@
$\mathcal{L}$ $=\alpha_{vae}\mathcal{L}_{vae}+\alpha_{pm}\mathcal{L}_{pm}+\alpha_{wm}\mathcal{L}_{wm}$ (8)
The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (237M parameters). A overview of the model architecture is in Fig. 6. To train our baseline models, we collect training data include $120$K episodes of demonstration from demonstration of privileged agents for imitation learning. Each episode has $2,048$ steps, which gives $240$M frames in total. To avoid the growing “compounding errors” (Ross and Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from dataset aggregation (DAgger) methods (Ross et al., 2011). We collected the trajectory of states by running a very noisy behavior policy, but used the privileged agent (100%) to label its action at each step (which is not actually executed). Additionally, to enhance the stability of auto-regression, we dropout and randomize $15\%$ of the input observations to the causal transformer during training (Details in Appendix A.2).
Figure 6: An overview of the model evaluated on the maze world tasks. Initially, a Variational Auto-Encoder (VAE) is trained to encode observed images into hidden vectors and vice versa, with its parameters subsequently frozen. The sequence of observations and actions is then utilized to train a causal model, specifically a transformer that employs only backward attention, using imitation learning and self-supervised learning techniques. For context-free and partial-context solutions, we intentionally mask some of the backward connections.
To investigate the impact of context length on performance, we introduce additional two variants that share the same structure as the small-sized causal transformer (with $26$M parameters), but have limited access to the contexts. The partial-context causal transformer has an attention window of 2, providing effective context length up to 9. The context-free transformer has an attention window of 1, providing effective context length of 1. The context-free and partial-context transformers are trained by further fine-tuning on the foundation of the full-context version (Appendix A.1).
@@ 3.3 Baseline Performances
Figure 7: Static assessment of the ability to imitate previleged agents ($\mathcal{L}_{pm}$) and predict the next frame ($\mathcal{L}_{wm}$) across each time step ($t$).
Static Evaluation. We first conduct static evaluation using off-policy data from an independently sampled validation set comprising $1275$ sequences (totaling $2.6$M frames). To analyze the impact of context length on performance, we assessed the losses of the world model and policy model for each individual time step ($\mathcal{L}_{wm}(t)$ and $\mathcal{L}_{pm}(t)$). In Figure 7, We compare the performance of the four variants, with several noteworthy observations: First, the ability to encode long contexts significantly influences performance, while the scale of the parameters shows no observable effect, which is also consistent with the conclusion of Meta-Language. Second, both world-modeling and policy-modeling improve with increased context length, though they exhibit some subtle differences. The capability of world-modeling improves rapidly within the first 20 steps, after which the rate of improvement slows. This could be because single-step world modeling is primarily impacted by the localized environment, explaining the rapid initial improvement with short contexts, but is less influenced by distant layouts, leading to slower improvements over longer contexts. In contrast, the improvement in policy modeling with increased context length is not monotonic, with the loss remaining relatively unchanged between $10$ and $500$ steps. The decision modeling loss is closely related to the reference policy, which follows an exploration-then-exploitation strategy. These results may indicate that imitating the exploration strategy is more challenging.
@@
Figure 8: Interactive evaluation results of the randomized, previleged and casual-modeling based agents, depicted by plotting accumulated rewards against simulated time steps. The step minor cost is set to 0. Each plot summarizes the results from static 64 tasks. The shaded area represents the 95% confidence interval of the expectation.
@@
Interactive Evaluation. It is more beneficial to investigate the online performance of the policy model and world model by directly interacting with the environment. We first directly evaluate the performance of the policy model by interacting with a set of evaluation tasks. The evaluation set includes $64$ tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$. We plot the average accumulated rewards and their confidence over time (t). For comparative analysis, we also include the performance of a random policy, where actions are randomly sampled regardless of the observations, as well as privileged agents with $p(\text{STM}\rightarrow\text{LTM})$ of $5\%$, $25\%$, $100\%$. The results (Figure 8) indicate that the model with full contexts outperforms the random policy, context-free agents, and privileged agents ($5\%$), but there is a significant gap compared to privileged agents with $25\%$ and $100\%$ LTM. However, when compared with partial-context causal transformers, the superiority of the full-context model is not obvious. Considering the conclusion in static evaluation, it is reasonable to consider that the performance of interactive evaluation is constrained by imitation learning itself. We believe it is beneficial to incorporate on-policy trajectories in the dataset in the future, either in reinforcement learning or DAgger (Ross et al., 2011).
(a) Maze $15\times 15$
(b) Maze $25\times 25$
(c) Maze $35\times 35$
Figure 9: Results regarding the in-context improvement of world modeling in Maze World. We examine the compound error associated with a 9-step auto-regressive future predicting process, using the learned world models that differ in the context length ($t\in\{1,100,1000,2000\}$). The figures presented compare different models to the context-free world model as the baseline. The horizontal axis measures the number of imagined steps, with each curve illustrating a distinct context length. The mean square error is averaged across 64 tasks, and the shaded area provides a 95% confidence interval for these measurements.
In-Context Improvement of World Modeling. Figure 9 explores the performance of world models in interactive evaluations. At steps $1$, $100$, $1000$, and $2000$ of each task, the world model is tasked with “imagining” $9$ steps into the future. For the smaller $15\times 15$ mazes, we observe a significant improvement in world modeling as context length increases from $1$ to $1000$, evidenced by a continuous decrease in prediction error. However, this improvement plateaus between $1000$ and $2000$ steps. With sufficient context, the cumulative error over 9 steps remains lower than that of an average guess, which is the performance of the context-free model. For larger mazes of $25\times 25$ and beyond, and when predicting more than 5 steps into the future, none of the trained models outperform an average guess. Additionally, the partial-context model does not enhance its predictive capabilities as context length increases, which aligns with expectations. The results indicate that the proposed model demonstrates a reasonable capacity to utilize contextual information for future predictions, yet there is considerable room for improvement, especially in solving larger-scale mazes and the efficient utilization of long-term dependencies. For further insights and cases, refer to Appendix A.4.
@@ 4.1 Meta-learning and GPICL
Meta-learning encompasses a wide range of methodologies, including gradient-based adaptation (Finn et al., 2017), attention-based adaptation (Mishra et al., 2018), and memory-based adaptation (Santoro et al., 2016; Duan et al., 2016; Wang et al., 2022; Lu et al., 2024). Both attention-based and memory-based adaptations are closely related to in-context learning, especially general-purpose ICL. Since the discovery that large-scale language models (LLMs) can perform few-shot learning (Ouyang et al., 2022), Transformer architectures and LLMs are increasingly becoming the state-of-the-art for meta-learning and in-context learning (ICL). This includes applications ranging from few-shot supervised learning (Min et al., 2021; Garg et al., 2022; Dai et al., 2023; Li et al., 2023), in-context reinforcement learning (Chen et al., 2021; Laskin et al., 2022; Lin et al., 2023), to multi-modal reasoning and embodied control (Reed et al., 2022; Szot et al., 2023). However, typical pre-training models only address relatively simple few-shot learning tasks, while more complex tasks still require task-specific or class-specific training. It is desirable to introduce general-purpose in-context learning (Kirsch et al., 2022; 2023) for generalizing across a wider variety of tasks, which necessitates benchmarks that can scale up in both quantity and diversity.
4.2 Relative Benchmarks
Meta-Learning Benchmarks. A variety of benchmarks have been utilized to evaluate meta-learning algorithms, encompassing domains such as image classification (LeCun et al., 1998), video games, and robot manipulation (Brockman et al., 2016). Many of these benchmarks were not initially designed for meta-learning but have been adapted to accommodate varying tasks through parametric adjustments in hyperparameters and by artificially hiding certain parameters from observation. This can be further advanced to include precedurely generated random targets of control (Finn et al., 2017; Mishra et al., 2018; Yu et al., 2020; Morad et al., 2023), video games of different difficulty level (Cobbe et al., 2019), geometry of locomotion (Najarro and Risi, 2020), rules of transition (Nikulin et al., 2023), and layouts of mazes (Mishra et al., 2018; Morad et al., 2023). However, some of these hidden configurations are relatively simple and fail to generate sufficient diversity, causing models to lean towards acquiring zero-shot and task identification abilities rather than in-context learning from scratch. For example, while many previous benchmarks include hiding the target of locomotion, an ideal learning algorithm should be able to locate the hidden target in at most three steps by comparing the rewards of walking in different directions, given that the location of the target is only in two dimensions. Another promising approach is to randomize labels (Kirsch et al., 2022; Wei et al., 2023), observations (Morad et al., 2023), and actions (Sinii et al., 2023), which tend to have higher dimensions. Based on the prior studies, it is plausible to infer that the diversity of tasks and the resulting generalization capabilities are related to the number of hyperparameters that can be varied. In contrast to the artificial randomization of rewards, observations, and targets, we believe that Meta-Language and MazeWorld offer environments that are inherently more complex and significant for benchmarking GPICL.
Language Modeling. Recently, emphasis has been placed on long-term dependency modeling in large language models (LLMs). Most of the newly proposed benchmarks support context lengths over 8K tokens (Rae et al., 2019; Tay et al., 2020; Bai et al., 2023; Shen et al., 2023; Song et al., 2024). However, there are several limitations when using current language-based benchmarks for in-context learning (ICL) investigation. First, some benchmarks focus on one-step classification with very long feature descriptions (Tay et al., 2020; Chalkidis et al., 2019; Bai et al., 2023), which is not suitable for examining the interactive performance of the model. Second, it is unclear how much long-term dependency actually exists in these long contexts (Chen et al., 2024). Even if some long-term dependencies are present, they may be weak and relatively simplistic (such as retrieval and copy tasks), and could be overshadowed by the powerful zero-shot generalization capabilities of the models. This overshadowing can obscure crucial information important for GPICL. In practice, it is often found that scaling up model parameters to enhance zero-shot generalization typically yields better results than extending the ICL horizon.
Visual Navigation. There are a few datasets featuring indoor and outdoor robot-environment interaction scenes, reconstructed from images taken in real-world settings (Song et al., 2017; Chang et al., 2017; Xia et al., 2018), or simulated environments (Szot et al., 2021). Currently, most of these environments primarily facilitate the learning of zero-shot generalization. A significant limitation of these datasets is the high data collection cost and the challenge of scaling up. Compared to these benchmarks, the proposed Maze World has a much larger sim-to-real gap, but offers greater diversity and complexity, and is much easier to scale up. Additionally, we suggest that reality-generated data could better serve as contexts in the GPICL stage, as illustrated in Figure 2(b), rather than during meta-training.
@@ 5 Conclusions
In this paper, we introduce two benchmarks specifically designed for GPICL. Our preliminary investigations of these benchmarks have yielded inspiring results, demonstrating that scaling laws might exist between context length and performances, and that the performance is not entirely dependent on the scaling of parameters.
@@
It is important to note that although we are seeing the potential of application of those datasets in areas including language modeling and indoor navigation, the synthetic data generated by those benchmarks are not yet validated to be ready for a direct help to application. This work can be expanded along two directions: 1. The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2. Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.
Appendix A Appendix
A.1 Model Description
Table 1: Comprehensive configurations of the baseline models
Model parameters layers $d_{model}$ $n_{heads}$ batch size (tokens)
Meta-Language Model (tiny) $303$K 6 64 4 $32$K
Meta-Language Model (small) $9.5$M 12 256 8 $32$K
Meta-Langauge Model (standard) $151$M 12 1,024 16 $64$K
Causal Transformer (small) $26$M 8 512 8 $32$K
Causal Transformer (standard) $237$M 12 1,280 16 $32$K
Figure 10: Structures of the encoder and decoder in the Variational Autoencoder
The comprehensive configurations of the baseline models are outlined in Table 1. Notably, within the Maze World framework, each temporal increment is encoded by a pair of tokens: an observation and its corresponding action. For the optimization of the Meta-Language and Maze World baseline models, we employed the Noam decay scheduler [Vaswani et al., 2017]. This scheduler peaks at a learning rate of $10^{-3}$ following an initial warm-up phase spanning $1,000$ steps.
For the image encoder (11M parameters) and image decoder (14M parameters), we employ convolutional and deconvolutional layers augmented with ResNet blocks [He et al., 2016]. The detailed structure of the VAE is depicted in Figure 10. Initially, we train the VAE in isolation by setting $\alpha_{wm}$ and $\alpha_{pm}$ in Equation 8 to 0. Subsequently, we freeze the parameters of the encoder and decoder and train the causal transformer only.
For partial-context and context-free causal transformers, we modify the attention window to 2 and 1, respectively, as shown in Figure 11. Since transformers can propagate information between layers, an attention window of 2 results in an effective context size of the number of layers plus 1.
(a) Full-context
(b) Partial-context
(c) Context-free
Figure 11: Variant attention masks of causal modeling
(a) Meta-Language Model
(b) World Modeling
(c) Policy Modeling
(d) World Modeling
(e) Policy Modeling
Figure 12: The training loss with respect to the steps of optimization. Notice that the partial-context and context-free causal transformers are initialized from the trained full-context transformers.
Figure 12 presents the training loss versus the number of iterations for the Meta-Language and Maze World baseline models. During training baselines for Meta-Language tasks, we observed that standard-sized models converge relatively slowly on the training set. To address this, we conducted additional warm-up epochs on a simpler pre-training dataset with $n=2,3$ before transitioning to the main pre-training dataset. During training the baselines for Maze World tasks, we found there was mutual interference between the world modeling loss ($\mathcal{L}_{wm}$) and policy modeling loss ($\mathcal{L}_{pm}$), especially in the initial stages. Our meta-training underwent several phases where we continuously adjusted the loss weights $\alpha_{wm}$ and $\alpha_{pm}$ in Equation 8. To further accelerate and stabilize the convergence of the world model, in addition to these losses, we introduced an auxiliary loss $\mathcal{L}_{z}=\sum_{t}||\hat{z}_{t}-z_{t}||^{2}$ with a weight of $\alpha_{z}=0.05$. During the final stage, the hyperparameters are adjusted to $\alpha_{pm}=0.25$ and $\alpha_{wm}=0.70$. However, we recommend starting with $\alpha_{wm}>0.90$ and $\alpha_{pm}<0.10$ as the policy model training loss is more volatile initially. The context-free and partial-context causal transformers are initialized from the full-context causal transformer, and the hyperparameters remain unchanged.
A.2 Data Collection Strategy for Maze World
For dataset collection and static evaluation of Maze World, inspired by dataset aggregation [Ross et al., 2011] and noise distillation [Zisman et al., 2023], we utilize distinct behavior and reference policies. The behavior policy is determined by previleged agent where $p(\text{STM}\rightarrow\text{LTM})$ randomly sampled between $[0,50\%]$. Additionally, with the probability of $\epsilon\in[0,80\%]$ the behavior action is determined by random decision. With the states generated by behavior policy, we use the reference policy with $p(\text{STM}\rightarrow\text{LTM})=100\%$ to label the action.
We’ve also randomly dropped 20% of the input observations ($z_{t}$) to enhance the robustness of auto-regression. Among these, 10% of the input observations are replaced with a special mask token with trainable embeddings, while the remaining 10% are perturbed with Gaussian noise $\mathcal{N}(0,1)$.
A.3 Additional Results in Meta-Language Model
(a) English vocabulary learning task.
(b) Mathematical operation task.
A.4 Additional Results in Maze World Tasks
(a) Random
(b) Context-free
(c) Partial-context
(d) Full-context
(e) Privileged Agent
Figure 14: Performances of different methods in 3 randomly picked $35\times 35$ tasks. Green trajectories indicate time steps $t$ close to 0, while red trajectories represent $t$ close to $T$. From left to right, the columns are generated by: (a) Random Policy (b) Context-free Model (c) Partial-context Model (d) Full-context Model (Small-sized) (c) Previleged agent with $p(\text{STM}\rightarrow\text{LTM})=100\%$.
In Figure 14, We display trajectories generated by different policies on three randomly selected $35\times 35$ tasks. Interestingly, we find that the main bottleneck of performance might arise from a lack of efficient exploration. Models with less context tend to explore less area. For instance, the random policy can only traverse a very small part of the maze. This observation could also explain the extraordinary challenge posed by the tasks: Traditional randomized exploration, a common technique in reinforcement learning, proves inadequate because it often revisits the same areas repeatedly without retaining information about which parts have been explored, thereby failing to cover broader regions effectively. Therefore, the agent must not only learn how to memorize paths and navigate but also how to explore more effectively, generating high-quality “training data” within its context to improve its future performance.
Figure 15: Cases regarding the in-context improvement of world modeling (with $15\times 15$ mazes, small-sized full-context causal transformer). Each clip predicts $9$ steps into the future, conditioned on a fixed sequence of actions, and based on varying lengths of context denoted by $t$. Each clip consists of two sections: the upper section displays the ground truth observations from the current step to $9$ steps ahead (from left to right), while the bottom section shows the predicted future states up to $9$ steps ahead.
In Figure 15, we illustrate a qualitative evaluation of the full-context causal transformer’s world modeling capabilities. Across the majority of instances investigated, a noticeable enhancement in future predicting is evident as the context length grows, suggesting that the model progressively adapts to the intricacies of the maze. This observation indicates the promising applicability of GPICL to the realm of adaptive embodied intelligence.
A.5 Additional Diversity in the Maze World
(a) 2D discrete
(b) SURVIVAL tasks
(c) narrow
(d) wide
(e) low
(f) high
(g) sparse
(h) dense
(i) close-sighted
(j) far-sighted
Figure 16: Additional variants by provided configurations in Maze World.
It is worth noting that Maze World is a lightweight and fast engine with an easy-to-use Python API based on GYM [Brockman et al., 2016] for simulating navigation in various mazes. It runs smoothly at over 20 frames per second on a standard CPU. In addition to the aforementioned details, Maze World incorporates various configurations and hyper-parameters that can be adjusted to create a wide range of tasks. For example, it offers different types of observation spaces and action spaces, including discrete and continuous actions, as well as 2D and 3D observations (Figure 16a). Beyond navigation tasks, we introduce more challenging survival tasks, where the rewards provided by the PNTs are static within an episode but initially unknown (can be negative, Figure 16b). The basic environments can also be modified in terms of the size of the basic cell (Figure 16c,16d), the height of the ceiling (Figure 16e,16f), the density of obstacles (Figure 16g,16h), and the range of view (Figure 16i,16j).
A.6 Details of Task Configurations
In Table 2, we list the detailed settings for the tasks used in this paper. The rewards for reaching the target are by default related to the size of the maze. Additionally, we note that the density of obstacles controls the number of loops in the environments. We recommend using a density between 0.15 and 0.40. A density over 0.50 results in mazes without any loops, significantly reducing the difficulty of traversing the maze (or performing exploration), as an agent can easily traverse the entire world by always going right (or left). We believe this might be a crucial reason why, in Morad et al. [2023], where the agent is only required to navigate to a single target, context-free models can outperform memory-augmented models.
Table 2: Setting of task configurations used in this paper
Meta-Langauge
Embedding size of the randomized langauge generator 32
Hidden size of the randomized language generator 64
Normalize factor for softmax of the randomized language generator ($\lambda$) 5
Number of the hidden parameters ($|\theta|$, $n=2$) $7$K
Number of the hidden parameters ($|\theta|$, $n=4$) $11$K
Number of the hidden parameters ($|\theta|$, $n=8$) $19$K
Maze Tasks
Density of obstacles $0.36$
Number of Potential Navigation Targets (PNTs) $10$
Resolution of observation $128\times 128$
Range of view 12.0 (default)
Size of basic cell 2.0 (default)
Rewards of arriving target ($15\times 15$) 0.57 (default)
Rewards of arriving target ($25\times 25$) 1.24 (default)
Rewards of arriving target ($35\times 35$) 2.06 (default)
Height of ceiling 3.2 (default)
Height of agents’ view points 1.6 (default)